Consider a pair of planets for which the distance between them is increased by a factor of three. Show what happens to the force of attraction between these two planets.
Answer #1

F = GMm/r^2


r' = 3r

so

F' = GMm/(3r)^2 = GMm/9r^2 = F/9

so Force is decreasedby 9 times